OpenStudy (anonymous):

What volume of O2 (g) is produced from 1.0 g BaO upon its decomposition to Ba(s) and O2(g) at STP? What volume of O2 (g) is produced from 1.0 g BaO upon its decomposition to Ba(s) and O2(g) at STP? @Chemistry

OpenStudy (anonymous):

0.0731L or 73.1 mL because n(BaO)= 1/153.3 = 0.006523 mole ratio = 2:1 therefore the moles of oxygen = 0.006523/2 = 0.00326 n=v/22.41 so v=n*22.41 = 0.00326*22.41 = 0.0731